OpenStudy (anonymous):

Where is this function concave up and concave down. x^2/(x-5)^2

OpenStudy (klimenkov):

\[\frac{x^2}{(x-5)^2}\]Can you take derivatives?

OpenStudy (anonymous):

yes. I know i have to take the second derivative, but i having trouble finding out where the 2nd derivative equals zero.

OpenStudy (anonymous):

I even put the function into a graphing calculator. i still cant get the right answer.

OpenStudy (klimenkov):

Can you type here what you've got?

OpenStudy (anonymous):

going to take me a while to get it, have to do chain rule and everything. '_' you can just post the answer if you want. >.> already been on this problem for like an hour

OpenStudy (klimenkov):

Ok. I will type an answer for you.

OpenStudy (anonymous):

this was a multiple part question, this is what i have so far. http://gyazo.com/003eb77ccf9977811aa71cf7f65fb35f

OpenStudy (anonymous):

all those answers are right. just need 7 and 8

OpenStudy (klimenkov):

\(f''(x)=\frac2{(-5 + x)^2} - \frac{8 x}{(-5 + x)^3} + \frac{6 x^2}{(-5 + x)^4}\) The horizontal asymtote isn't right.

OpenStudy (klimenkov):

Oh. Sorry. I must check it again!

OpenStudy (klimenkov):

Sorry, I'm not very good in English. Am I right?